OpenStudy (anonymous):

I am confused about a part of my Algebra work. Please help me out if you can thank you... Now that you have helped Bernard design the pool, he presents you with a situation. There are x number of communities that will be designing homes with the pools. The number of homes within each community is six more than the number of communities. The number of hours that will be required to build each home is double the number of homes in each community. I have the answers for this but the next part is what I don't get...

OpenStudy (anonymous):

Using the three expressions, interpret the meaning of each grouping of factors as a single unit and also simplify each expression by finding their product.

OpenStudy (anonymous):

There are x number of communities that will be designing homes with the pools. so no. of communities = x The number of homes within each community is six more than the number of communities. so the total number of homes = x(x+6) The number of hours that will be required to build each home is double the number of homes in each community. total hours= 2(x+6)

OpenStudy (anonymous):

Ok I also got confused @triciaal can you help?

OpenStudy (anonymous):

Okay

OpenStudy (studygurl14):

Do you still need help @equestrianlife @sangya21 ?

OpenStudy (anonymous):

Yups @StudyGurl14

OpenStudy (triciaal):

from above so the total number of homes = x(x+6) The number of hours that will be required to build each home is double the number of homes in each community. total hours= 2(x+6) should be 2{x(x+6)} next interpret the meaning of each grouping of factors as a single unit I think this is just restating what is given if the number of communities = x then the number of homes is a function of x such that it is x^2 + 6 x and the time required to build each home 2x^2 + 12 x as a single unit when there is only one community, x = 1 then there will be 7 homes and it would take 14 hours to build That's just what I got from it. You said you have the answers. Does my view help?